# Inducing Causal World Models in LLMs for Zero-Shot Physical Reasoning

### v1

@@ I Introduction
The last few years have witnessed a paradigm shift in artificial intelligence, driven by the advent of Large Language Models (LLMs). Foundational models like BERT [1] and the GPT series [2, 3] have demonstrated remarkable, often emergent, capabilities across a vast spectrum of natural language tasks. Their proficiency is not limited to text; recent multimodal models such as Gemini [4] and Flamingo [5] have extended this power to encompass vision and other modalities, achieving impressive results in areas like visual dialog [6] and video question answering [7]. The capabilities of these models are so extensive that they are often evaluated on large, complex benchmark suites [8] and are increasingly being integrated with external tools to further enhance their functionality [9].
@@
Despite their success in processing and generating human-like language, these models harbor a fundamental limitation: they operate primarily by learning statistical correlations from data, lacking a deep, grounded understanding of the world. This gap becomes particularly apparent when LLMs are confronted with tasks requiring intuitive physical reasoning [10]. While they can often recite physical laws from their training data, they frequently fail to apply these principles to novel scenarios, a limitation that has been extensively documented [11]. Advanced prompting strategies such as Chain-of-Thought [12] and Tree-of-Thoughts [13] have been proposed to elicit more structured reasoning. However, these methods are essentially techniques for organizing the model’s existing knowledge rather than instilling a genuine, predictive understanding of physical dynamics [14]. This deficiency hinders the deployment of LLMs in high-stakes, real-world applications like robotics and autonomous systems, where the ability to predict the consequences of actions is paramount [15, 16]. The challenge is not merely about recognizing objects, but about understanding their interactions, a problem also explored in audio-visual event analysis [17].
@@
The core of the issue lies in the distinction between correlation and causation. Current models excel at the former, but true physical intelligence requires the latter. This has led to a growing interest in building models that are more physically grounded [18, 19]. Efforts in this direction include creating embodied multimodal models [20], developing new benchmarks for physical and event-based reasoning [21, 22, 23], and exploring novel architectures for efficient video analysis [24]. Concurrently, research in other domains has focused on interpreting the physical world through diverse signals, from WiFi-based gesture and activity recognition [25, 26] to emotion sensing [Gu2023WiFE] and even keystroke eavesdropping [27], demonstrating the broad effort to connect AI to physical phenomena. However, these specialized systems often focus on classification rather than building a general, predictive model of the world. The development of such general models is computationally intensive, driving research into efficient methods like federated learning [28] and model compression [29].
@@ II-B World Models and Physical Prediction
The concept of a ”world model” [38], an internal, predictive model of an environment, originated largely in the context of reinforcement learning and robotics [30]. Models like DreamerV3 [39] learn directly from sensory inputs to create a latent space where planning can occur, enabling agents to master diverse and complex tasks. In parallel, the computer vision community has developed models for explicit physical prediction from visual data. Physics-informed neural networks (PINNs) embed differential equations directly into the learning process [40]. Other work has focused on creating challenging benchmarks like CLEVRER [21] and Physion [23] to evaluate models on their ability to predict the outcomes of physical events. Recently, generative models like GENIE [41] have shown the ability to create entire interactive environments from a single image, demonstrating a powerful form of implicit world modeling. Our work draws inspiration from these approaches but differs in a key aspect: instead of learning from low-level sensory data or within a constrained simulation, we aim to induce a general, causal world model within a pre-trained LLM using a combination of textual descriptions and video observations. This leverages the vast knowledge already present in the LLM while grounding it in causal physical principles. The development of such complex models relies on advances in underlying neural architectures, such as Mamba [42], and efficient learning schemes across heterogeneous networks [43].
@@ II-C Causality and Multimodal Grounding
A central thesis of our work is that true physical reasoning requires a causal understanding. The field of causal inference provides a formal language for this, but applying it to high-dimensional data like text and video is a significant challenge [33]. Some work has explored using counterfactuals for data augmentation [44] or learning causal relationships in a compositional manner [34]. In the multimodal domain, significant effort has been directed at grounding language in perception. This includes pre-training on large-scale video-language datasets [45, 46], developing neuro-symbolic architectures that combine perception with explicit reasoning [47], and creating embodied agents that learn from interaction [48]. The challenge of grounding is also present in other sensing modalities, where researchers aim to prevent attacks on physical-layer authentication [49] or build robust activity recognition systems that are immune to interference [50, 51]. These efforts to build reliable systems in the face of environmental noise or adversarial perturbations [42] share a common goal with our work: to move beyond superficial correlations. While a system might learn to associate a specific WiFi signal pattern with an activity [52], it does not understand the underlying causal physics of human movement. Our work directly tackles this by using causal interventions as a supervisory signal, forcing the model to learn not just *what* happens, but *why* it happens. This is analogous to efforts in other fields to learn from heterogeneous data sources [53] or perform commonsense reasoning over knowledge graphs [54].

### v2

@@ I Introduction
The last few years have witnessed a paradigm shift in artificial intelligence, driven by the advent of Large Language Models (LLMs). Foundational models like BERT [1] and the GPT series [2, 3] have demonstrated remarkable, often emergent, capabilities across a vast spectrum of natural language tasks. Their proficiency is not limited to text; recent multimodal models such as Gemini [4, 5] and Flamingo [6] have extended this power to encompass vision and other modalities, achieving impressive results in areas like visual dialog [7] and video question answering [8]. The capabilities of these models are so extensive that they are often evaluated on large, complex benchmark suites [9] and are increasingly being integrated with external tools to further enhance their functionality [10].
@@
Despite their success in processing and generating human-like language, these models harbor a fundamental limitation: they operate primarily by learning statistical correlations from data, lacking a deep, grounded understanding of the world.[11]This gap becomes particularly apparent when LLMs are confronted with tasks requiring intuitive physical reasoning [12]. While they can often recite physical laws from their training data, they frequently fail to apply these principles to novel scenarios, a limitation that has been extensively documented [13]. Advanced prompting strategies such as Chain-of-Thought [14] and Tree-of-Thoughts [15] have been proposed to elicit more structured reasoning. However, these methods are essentially techniques for organizing the model’s existing knowledge rather than instilling a genuine, predictive understanding of physical dynamics [16]. This deficiency hinders the deployment of LLMs in high-stakes, real-world applications like robotics and autonomous systems, where the ability to predict the consequences of actions is paramount [17, 18]. The challenge is not merely about recognizing objects, but about understanding their interactions, a problem also explored in audio-visual event analysis [19].
@@
The core of the issue lies in the distinction between correlation and causation. Current models excel at the former, but true physical intelligence requires the latter. This has led to a growing interest in building models that are more physically grounded [20, 21]. Efforts in this direction include creating embodied multimodal models [22], developing new benchmarks for physical and event-based reasoning [23, 24, 25], and exploring novel architectures for efficient video analysis, such as unified networks for temporal filtering [26]. This is part of a broader trend in creating generative interactive environments from visual data [27]. Concurrently, research in other domains has focused on interpreting the physical world through diverse non-visual signals. WiFi-based sensing, for example, has proven to be a versatile modality. Foundational work established robust human activity recognition by developing methods to mitigate co-channel interference [28], a crucial step for any real-world interactive system [29]. This reliability is paramount for embodied agents that must ground language in robotic affordances [17]. Building on this, more complex applications like gesture recognition have emerged, where systems must not only be accurate but also handle unknown, open-set gestures to be practical in the real world [30]. The application space for such sensing has expanded dramatically, a trend seen across many forms of human-AI interaction [31], moving into contactless healthcare with systems that can capture pulmonary function [32]. This extends beyond WiFi, with other RF technologies like RFID-based systems also being explored for identifying fine-grained activities such as writing [5]. Such sensing paradigms enable nuanced tasks that require a form of machine commonsense to interpret [33], including unobtrusive emotion recognition [34]. However, the increasing sensitivity of data collected by such pervasive sensing also raises significant security concerns, leading to research into threats like keystroke eavesdropping from commodity devices [35]. These specialized systems, while powerful, often focus on classification rather than building a general, predictive model of the world. The development of such general models is computationally intensive, driving research into distributed learning schemes like federated learning with experience-driven model migration [36]. This is complemented by a focus on algorithmic efficiency through novel architectures [37] and advanced techniques like multi-objective model compression [38].
@@ II-B World Models and Physical Prediction
The concept of a ”world model” [47], an internal, predictive model of an environment, originated largely in the context of reinforcement learning and robotics [39]. Models like DreamerV3 [40] learn directly from sensory inputs to create a latent space where planning can occur, enabling agents to master diverse and complex tasks from rich data. This includes understanding and predicting fine-grained micro-actions, a challenge addressed by dedicated benchmarking efforts [48]. In parallel, the computer vision community has developed models for explicit physical prediction from visual data. Physics-informed neural networks (PINNs) embed differential equations directly into the learning process [49]. Other work has focused on creating challenging benchmarks like CLEVRER [23] and Physion [25] to evaluate models on their ability to predict the outcomes of physical events. Recently, generative models like GENIE [27] have shown the ability to create entire interactive environments from a single image, demonstrating a powerful form of implicit world modeling. Our work draws inspiration from these approaches but differs in a key aspect: instead of learning from low-level sensory data or within a constrained simulation, we aim to induce a general, causal world model within a pre-trained LLM using a combination of textual descriptions and video observations. This leverages the vast knowledge already present in the LLM while grounding it in causal physical principles. The development of such complex models relies on advances in underlying neural architectures, such as Mamba [37], and efficient learning schemes across heterogeneous networks [36].
@@ II-C Causality and Multimodal Grounding
A central thesis of our work is that true physical reasoning requires a causal understanding. The field of causal inference provides a formal language for this, but applying it to high-dimensional data like text and video is a significant challenge [42]. Some work has explored using counterfactuals for data augmentation [50] or learning causal relationships in a compositional manner [43]. In the multimodal domain, significant effort has been directed at grounding language in perception. This includes pre-training on large-scale video-language datasets [51, 52], developing neuro-symbolic architectures that combine perception with explicit reasoning [53], and creating embodied agents that learn from interaction [54]. The challenge of grounding is also present in other sensing modalities, where researchers aim to prevent attacks on physical-layer authentication [55]. This focus on security and reliability is paramount for embodied agents that must function in noisy, real-world conditions [17]. In a similar vein, others seek to build robust activity recognition systems by developing anti-interference techniques based on phase components [56]. The ability to extract a clean signal from noise is a general problem, with solutions often involving novel neural module communication strategies [57]. The successful application of these robust techniques enables more advanced, target-oriented sensing in cluttered environments for applications like healthcare [58]. Interpreting these signals to understand high-level human states, such as emotions, requires a form of intuitive physics or theory of mind [59]. This is exemplified by systems that fuse WiFi and vision for unobtrusive emotion recognition [34]. Our work directly tackles this by using causal interventions as a supervisory signal, forcing the model to learn not just *what* happens, but *why* it happens. This is analogous to efforts in other fields to learn from heterogeneous data sources to a more reliable representation [60] or perform commonsense reasoning over knowledge graphs [33].